-4x+5y=6
3x+2y=7
x=? y=?

Answer by checkley77(12844)

-4x+5y=6 mulyiply this equation by 3
-12x+15y=18
-------------------
3x+2y=7 multiply this equation by 4
12x+8y=28
now add them:
----------------
-12x+15y=18
12x+8y=28
----------------------
23y=46
y=46/23
y=2 answer.
-4x+5*2=6
-4x+10=6
-4x=6-10
x=-4/-4
x=1 answer.